OpenStudy (anonymous):

Can anyone help? 2*log(x-5)=0 solve for x

OpenStudy (anonymous):

2 times what = 0?

OpenStudy (anonymous):

\[2\times \log_{10}(x-5)=0 \]

OpenStudy (anonymous):

I should have said divide both sides by 2 to start out with. Then you will have log (x-5) = 0. So what value do we need for x.